A motor cyclist travelling at 12m/s decelerates at 3m/s2.

a) How long does it take?

b) how far does he travel in coming to rest?


1
Expert's answer
2022-06-01T12:48:54-0400

(a)

"t=\frac{\Delta v}{a}=\frac{-12\;\rm m/s}{-3\:\rm m/s^2}=4\:\rm s"

(b)

"d=v_0t+at^2/2\\
=12*4+(-3)*4^2/2=24\:\rm m"